# Misquoted No More: Securely Extracting F Programs with IO

### v1

@@ Abstract.
We apply this general idea to build SEIO, a framework for extracting shallowly embedded F programs with IO to a deeply embedded $\lambda$-calculus while providing formal secure compilation guarantees. Using two cross-language logical relations, we devise a machine-checked proof in F that SEIO guarantees Robust Relational Hyperproperty Preservation (RrHP), a very strong secure compilation criterion that implies full abstraction as well as preservation of trace properties and hyperproperties against arbitrary adversarial contexts. This goes beyond the state of the art in verified and certifying extraction, which so far has focused on correctness rather than security.
@@ 1. Introduction
Artifact. This paper comes with an artifact in F that contains SEIO and a complete machine-checked proof that it satisfies RrHP. The proofs of some of the low-level lemmas we stated in this formalization were completed by Claude Opus 4.6, which is extremely good at doing F proofs with minimal human guidance (Swamy, 2026). The artifact includes over 900 top-level definitions (including type and function definitions, lemmas, and test programs) and for around 75 of them the actual definition was produced by Claude, usually lemmas for which we had written the statement and Claude produced the complete proof including by adding additional lemmas in a vibe proving session that took a couple of days and would cost a Claude user less than 15\$. The largest such proofs are over 200 lines long, look natural to us, and would have taken us much longer to write. The artifact is available at: https://github.com/andricicezar/fstar-io/tree/icfp26/seiostar
@@ 2.3. Formally secure extraction from IO to $\lambda_{io}$
We can use the trace-producing semantics of IO to verify the wrapper, since it allows us to write postconditions over a final result and a local trace of events. We can verify the wrapper to satisfy a safety property in arbitrary contexts: which states that for any file, task and agent, if the wrapper runs successfully (the Inl case), looking at the trace of events, the contents last read from the file corresponds (according to the validate function) to performing the task on the contents first read from the file.
@@ 5.2. Robust Relational Hyperproperty Preservation (RrHP)
We show that SEIO satisfies Robust Relational Hyperproperty Preservation (RrHP), the strongest secure compilation criterion of Abate et al. (2019). RrHP not only implies full abstraction (i.e., preserving observational equivalence), but also ensures the preservation of trace properties and hyperproperties (like noninterference) against adversarial contexts.
@@
The statement of RrHP uses an equivalence, so we prove each direction separately. In both cases, to instantiate the existential for source contexts, we use a back-translation function (denoted by $C^{T}\hskip-3.50006pt\uparrow$) from target contexts to source contexts. Since target contexts are syntactically well-typed, back-translation is defined recursively on the typing relation of $\lambda_{io}$. We prove in F that the source context obtained by back-translation is related to the original target context with respect to both our logical relations. These results are another instance of the fundamental property of logical relations for our two relations, so the proofs again rely on compatibility lemmas discussed in § 4.6.
To prove the right-to-left direction of RrHP ($\Rightarrow$), we use the target-to-source logical relation that relates each target behavior to a source behavior (§ 4.5).
@@
$\forall I^{S}.\ \forall C^{T}.\ \exists C^{S}.\ \forall P.\ \forall t.\ C^{T}[P\hskip-3.50006pt\downarrow]\leadsto_{T}t\Rightarrow C^{S}[P]\leadsto_{S}t$
@@
We back-translate the target context to get a source context to instantiate the existential. Since compilation is verified and $P$ is a value, we know that $P$ valid_contains $P\hskip-3.50006pt\downarrow$. Since back-translation is verified and $C^{T}$ is a value, we know that $C^{T}\hskip-3.50006pt\uparrow$ valid_contains $C^{T}$. Since our programs are functions, it follows that $(P~C^{T}\hskip-3.50006pt\uparrow)$ valid_superset_prod (EApp $P\hskip-3.50006pt\downarrow C^{T}$). Since linking is function application, this is all we need. ∎
This direction of RrHP already implies Robust Trace Property Preservation (RrTP), which is a weaker criterion that RrHP, but still one of the strongest secure compilation criteria of Abate et al. (2019). They show that RrTP implies full abstraction, and also ensures the preservation of trace properties and subset-closed hyperproperties against adversarial contexts.
@@
$\forall I^{S}.\ \forall C^{T}.\ \exists C^{S}.\ \forall P.\ \forall t.\ C^{S}[P]\leadsto_{S}t\Rightarrow C^{T}[P\hskip-3.50006pt\downarrow]\leadsto_{T}t$

### v2

@@ Abstract.
We apply this general idea to build SEIO, a framework for extracting shallowly embedded F programs with IO to a deeply embedded $\lambda$-calculus while providing formal secure compilation guarantees. Using two cross-language logical relations, we devise a machine-checked proof in F that SEIO guarantees Robust Relational Hyperproperty Preservation (RrHP), a very strong secure compilation criterion that implies full abstraction as well as preservation of trace properties and hyperproperties against arbitrary linked adversarial code. This goes beyond the state of the art in verified and certifying extraction, which so far has focused on correctness rather than security.
@@ 1. Introduction
Artifact. This paper comes with an artifact in F that contains SEIO and a complete machine-checked proof that it satisfies RrHP. The proofs of some of the low-level lemmas we stated in this formalization were completed by Claude Opus 4.6, which is extremely good at doing F proofs with minimal human guidance (Swamy, 2026). The artifact includes over 900 top-level definitions (including type and function definitions, lemmas, and test programs) and for around 75 of them the actual definition was produced by Claude, usually lemmas for which we had written the statement and Claude produced the complete proof including by adding additional lemmas. The largest such proofs are over 200 lines long, look natural to us, and would have taken us much longer to write. The artifact is available at: https://github.com/andricicezar/fstar-io/tree/icfp26/seiostar
@@ 2.3. Formally secure extraction from IO to $\lambda_{io}$
We use the trace-producing semantics of IO to verify the wrapper, since it allows us to write postconditions over a final result and a local trace of events. We verify the wrapper to satisfy a safety property in arbitrary contexts: which states that for any file, task and agent, if the wrapper runs successfully (the Inl case), looking at the trace of events, the contents last read from the file corresponds (according to the validate function) to performing the task on the contents first read from the file.
@@ 5.2. Robust Relational Hyperproperty Preservation (RrHP)
We show in F that SEIO satisfies Robust Relational Hyperproperty Preservation (RrHP), the strongest secure compilation criterion of Abate et al. (2019). RrHP not only implies full abstraction (i.e., preserving observational equivalence), but also ensures the preservation of trace properties and hyperproperties (like noninterference) against adversarial contexts.
@@
The statement of RrHP includes an existential quantifier, $\exists C^{S}$, which we instantiate using a back-translation function (denoted by $C^{T}\hskip-3.50006pt\uparrow$) from target contexts to source contexts. Since target contexts are syntactically well-typed, back-translation is defined recursively on the typing derivations of $\lambda_{io}$. The body of RrHP is an equivalence, so we prove each direction separately by showing that the source context obtained by back-translation is related to the original target context with respect to one of our two logical relations. These results are another instance of the fundamental property of logical relations for our two relations, so the proofs again rely on compatibility lemmas from § 4.6.
To prove the right-to-left direction of RrHP ($\Rightarrow$), we use the logical relation that relates each target behavior to a source behavior (§ 4.5).
@@
$\forall I^{S}.\ \forall C^{T}.\ \forall P.\ \forall t.\ C^{T}[P\hskip-3.50006pt\downarrow]\leadsto_{T}t\Rightarrow C^{T}\hskip-3.50006pt\uparrow[P]\leadsto_{S}t$
@@
Since compilation is verified and $P$ is a value, we know that $P$ valid_contains $P\hskip-3.50006pt\downarrow$. Since back-translation is verified and $C^{T}$ is a value, we know that $C^{T}\hskip-3.50006pt\uparrow$ valid_contains $C^{T}$. Since our programs are functions, it follows that $(P~C^{T}\hskip-3.50006pt\uparrow)$ valid_superset_prod (EApp $P\hskip-3.50006pt\downarrow C^{T}$). Since linking is function application, this is all we need. ∎
This RrHP direction already implies Robust Relational Subset-Closed Hyperproperty Preservation (RrSCHP), which is a weaker criterion than RrHP, but still one of the strongest secure compilation criteria of Abate et al. (2019). They show that RrSCHP implies full abstraction, and also ensures the preservation of trace properties and subset-closed hyperproperties against adversarial contexts.
@@
$\forall I^{S}.\ \forall C^{T}.\ \forall P.\ \forall t.\ C^{S}\hskip-3.50006pt\uparrow[P]\leadsto_{S}t\Rightarrow C^{T}[P\hskip-3.50006pt\downarrow]\leadsto_{T}t$